# Chlorophyll-a Mapping and Prediction in the Mar Menor Lagoon Using C2RCC-Processed Sentinel 2 Imagery

### v1

@@ III-C Buoy data
The data gathered comes from two complementary sources [53] and [54]. The latter is no longer available due to the end of the project related to that website. For simplicity, the initial data source will be referred to as Polytechnic University of Cartagena (UPCT), and the subsequent one as Murcian Institute of Agricultural and Environmental Research and Development (IMIDA). It is hypothesized that both sources are located in the same locations; however, the metadata and information in the data sources do not specify if the buoys are the same. The UPCT dataset encompasses the period from 2017-05-19 to 2024-07-18. Buoys measure at increasing depths, ranging from 0.5 meters to 5.0 meters. In areas where it is feasible to descend to these depths, measurements are typically taken in steps of 0.5 meters. The IMIDA dataset ranges form 2016-06-29 to 2023-08-24 and measurements are taken at surface level (0.0 meters) and then at 1, 2, 3, 4 and 5 meters. In both cases, data is retrieved approximately once a week, despite variations over time and across seasons.
An analysis comparing both datasets at each depth reveals that, in most cases, the differences between the measurements are not significant. Therefore, this merged data goes from 2016-06-29 to 2024-07-18, close to a decade-long dataset.
@@
III-D Satellite imagery
Sentinel 2 imagery is accessible to the public via the Copernicus Browser [55]. Within this interface, users can manually define a bounding box to delineate the area of interest, specify a date range, select the desired level of processing, choose the data source and download SAFE files. These SAFE files correspond to a large area, particularly a tile, and contains an image file for each spectral band, along with metadata. A different approach that is considerably more suitable for the downloading of multiple files is through the Copernicus API, a method that permits the automation of the process. In our case, we use a Python script to achieve this objective.
This script was developed to access the Copernicus Data Space Ecosystem repository through the boto3 [56] interface, The dates of interest and the Sentinel 2 tile identifier corresponding to the study area (Tile ID: 30SXG) were specified. The script dynamically constructs the search prefixes based on the date, following the hierarchical structure of the repository. The specific designation for this satellite is ”Sentinel 2/MSI/L1C_N0500/<<year>>/<<month>>/<<day>>” or ”Sentinel 2/MSI/L1C/<<year>>/<<month>>/<<day>>”, depending on the date. For each shot, the product in SAFE format containing the required Tile ID is selected, and all spectral bands are downloaded together with the associated metadata files to a local directory. By downloading L1C products as files in SAFE format, we can process them later on with SNAP.
@@ IV Methodology
IV-A Buoy Data preprocessing
@@
IV-B Satellite data processing
The SAFE files are processed with SNAP. In order to apply the atmospheric correction algorithms from the C2RCC project this process is performed:
@@ IV-B Satellite data processing
    Resample the image to 10 m with the “Nearest” method.
@@
    Crop the subset of the image with the geocoordinates 37.82, -0.867, 37.62, -0.7 for north latitude, west longitude, south latitude, and east latitude bounds, respectively.
@@
    Apply de C2RCC processing for S2-SMI, selecting as outputs TOA reflectantes (uncorrected), normalized water leaving reflectances (named rhown) and atmospherically corrected angular dependent reflectances (rhow) [57].
@@
Once the flux is defined, it is possible to automate the process with the SNAP functionality Graph Processing Framework [58]. This framework is used with XML files that define the processing workflow described above, which is further customized through a properties file to manage the parameters of each step. To fully automate the process, a bash script is then employed. The script iterates through all the dates of interest, applies the workflow specified in the XML file, and saves the resulting outputs as TIFF files. There are three iterations for each date to obtain the image processed with C2RCC, C2X and C2X-Complex.
The TIFF files were processed using Python with the aim of aggregating pixels within windows of different sizes to assess which aggregation scale yields the most reliable results. For each of the C2RCC neural networks, five window sizes were tested: 1×1, 3×3, 5×5, 9×9, and 15×15 pixels. In practice, this means that reflectance values at the buoy locations were derived by calculating the mean of the surrounding pixels rather than relying solely on the exact pixel corresponding to the buoy position. This approach was adopted because the processed images often exhibited substantial pixel-to-pixel variation even within very small areas. By averaging across a window, local fluctuations are smoothed, providing a more representative estimate of the reflectance at a given location. The outputs of this procedure were stored as CSV files, with one file generated for each combination of neural network and window size.
IV-B1 Band combinations
Following the ideas from related works previously mentioned, along with more options provided by [59], many band combinations were explored as inputs for the models. In preliminary tests, formulas from [31], [25] and [33] where considered, but were finally discarded due to not improving the models. Therefore, the combinations used are:
@@ IV-B1 Band combinations
IV-C Datasets
With the buoy and satellite data prepared, the next step is combine them into merged datasets, and obtain a final dataset for each group of depth, processing, and window aggregation. Additionally, C2RCC, C2X and C2X-Complex add another layer of dimensionality because they derived into two versions: atmospherically corrected and normalized reflectances, named rhow y rhown, respectively; whereas TOA reflectances are obtained from one of the previous, and are a single set of bands since they have not undergone any processing. The merging of buoy and satellite data is performed exclusively when both measurement sources are collected on the same date. The proposal in [27] to include more observations by using satellite images taken one or two days before or after the chlorophyll measurement was considered. However, since chlorophyll is a highly variable parameter, it was decided to select only exact matches.
@@ IV-C Datasets
V Experiments
The experiments conducted are described in this section. The scripts are available in the Github repository https://github.com/Antonio-MI/mar-menor-chl. The experiments were run on a system with Ubuntu 22.04.2 LTS, 16 GB of RAM, and an AMD Ryzen 5 5600G processor. The code was implemented using Python 3.8.5 and SNAP processing with SNAP 11.0.
@@ V Experiments
The preliminary training stage began with setting the model parameters to their default values, removing missing entries, and training with a 5-fold stratified cross-validation scheme. Stratification was guided by an auxiliary variable, “High_Chl”, in order to enforce a uniform distribution of samples with high Chl-a concentrations, since these were scarce in the dataset and a purely random split could otherwise lead to poor performance. Using this cross-validation setup (described in more detail in the final training phase), the top 10 performing datasets were selected for each processing method. This was based only on the highest R2 and RMSE values. This means that, across all the algorithms, the best one determines whether a dataset is selected or not. The best results in this stage range from an $R^{2}$ of approximately 0.8 in the superficial layer to 0.6 in the deepest one.
After preliminary training, hyperparameter optimization was performed using the Optuna framework [60], which allows to create a “study” instance to optimize an objective function for a particular metric specified by the user. The framework also allows users to select a range of possible values for the hyperparameters, depending on whether the values are integers, floats, or categorical. The output of this process is a file for each depth containing the optimal parameters for each dataset and model pair.
Once the hyperparameters had been optimized for each dataset–defined by the target Chl-a depth, the processing method, and the reflectance aggregation window–the final training phase was carried out for each depth. Models were initialized the hyperparameters selected with Optuna. For each dataset the data were first split into training (75%) and test (25%) subsets. This scheme used the same training partition and folds as in the selection stage to ensure that data leakage was avoided. The split was stratified using an auxiliary high-chlorophyll class (Chl>>5, quantile 90%, named “High_Chl”) defined to ensure a balanced representation of scarce high-concentration samples. The input variables (X) and the target (y) were then separated, making sure that neither the target nor the high-chlorophyll indicator were included in X to prevent data leakage. Model evaluation was conducted within a 5-fold cross-validation scheme stratified by the high-chlorophyll class. For each model, predictions were generated on the validation folds while the corresponding true values and indices were stored, enabling later construction of the ensemble. Predictions for the test set were also saved at each fold, contributing with equal weight (1/5) to the final test estimate. For distance-based models, data were scaled before training and predictions were back-transformed, whereas tree-based models were trained directly. In all cases, $R^{2}$ and RMSE were computed on the validation folds and stored for later aggregation.
The ensemble was evaluated in an additional loop of 5 folds: for each fold, the meta-model was trained on predictions from the remaining four folds and validated on the held-out fold. This procedure provided five independent validation estimates, consistent with the evaluation of the base models.
Finally, metrics were also computed on the test set for each model, and all results were stored in dictionaries and serialized using Pickle for reproducibility.
V-A Results
@@
Model performance was evaluated using two complementary metrics: the coefficient of determination ($R^{2}$) and the root mean squared error (RMSE). $R^{2}$, computed with the r2_score function from the scikit-learn library [61], quantifies the proportion of variance in the observed values that is explained by the model, with values closer to 1 indicating a better fit. RMSE measures the average magnitude of the prediction errors, expressed in the same units as the target variable, and lower values denote higher accuracy. Formally, given observed values $y_{i}$, predicted values $\hat{y}_{i}$, and their mean $\bar{y}$, the metrics are defined as:
@@
Since models and datasets are grouped for each depth, result are also shown by depth. To ensure conciseness, we only report the test performance metrics, but train metrics are also available in the Supplementary Material. These metrics provide the most relevant assessment of the models’ predictive ability.
Each table consists of ten rows, one for each of the datasets, determined by type of processing, window aggregation and depth, and a column for each model.
Tables IV and V show results for the most superficial layer. The best-performing dataset was “C2X-Complex_rhow_9x9”, which achieved an $R^{2}$ of 0.89 with both the XGB algorithm and the ensemble. The corresponding RMSE was 1.41 mg/m3 for XGB and slightly higher for the ensemble (1.43 mg/m3 ). CatBoost also yielded competitive results, although slightly below those of XGB and the ensemble. The second-best performance was obtained with the same dataset but using the normalized reflectances rhown. Overall, datasets with larger aggregation windows provided better results; indeed, none of the top 10 configurations employed 1×1 or 3×3 windows. Regarding the processing approach, C2X-Complex consistently outperformed the alternatives, followed by C2RCC.
@@
TABLE IV: Test $R^{2}$ for depth 0-1 meters. The best model for each scenario is highlighted in bold.
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
C2X-Complex_rhow_9x9_depth_in_0_1 0.88 0.68 0.89 0.76 0.87 0.65 0.78 0.82 0.74 0.89
C2X-Complex_rhown_9x9_depth_in_0_1 0.86 0.70 0.86 0.78 0.85 0.65 0.82 0.80 0.74 0.84
C2RCC_rhow_5x5_depth_in_0_1 0.80 0.64 0.82 0.85 0.76 0.43 0.79 0.75 0.71 0.81
C2X-Complex_rhow_15x15_depth_in_0_1 0.82 0.68 0.84 0.85 0.81 0.68 0.78 0.79 0.72 0.81
TOA_15x15_depth_in_0_1 0.68 0.56 0.85 0.85 0.61 0.51 0.69 0.49 0.34 0.61
C2RCC_rhow_15x15_depth_in_0_1 0.77 0.58 0.84 0.84 0.71 0.61 0.77 0.76 0.70 0.78
C2RCC_rhown_5x5_depth_in_0_1 0.82 0.64 0.84 0.82 0.80 0.50 0.81 0.78 0.71 0.84
C2X-Complex_rhow_5x5_depth_in_0_1 0.84 0.53 0.76 0.77 0.84 -0.11 0.74 0.80 0.72 0.84
C2RCC_rhow_9x9_depth_in_0_1 0.81 0.67 0.80 0.79 0.80 0.65 0.74 0.76 0.72 0.83
C2X-Complex_rhown_15x15_depth_in_0_1 0.77 0.67 0.79 0.81 0.73 0.64 0.73 0.73 0.71 0.77
@@
TABLE V: Test RMSE (mg/m3) for depth 0-1 meters. The best model for each scenario is highlighted in bold.
Model CAT ELN ENS KNN LBM LR MLP RF SVR XGB
C2X-Complex_rhow_9x9_depth_in_0_1 1.51 2.45 1.43 2.10 1.58 2.56 2.00 1.82 2.18 1.41
C2X-Complex_rhown_9x9_depth_in_0_1 1.63 2.35 1.62 2.00 1.67 2.56 1.81 1.93 2.19 1.70
TOA_15x15_depth_in_0_1 2.42 2.82 1.63 1.66 2.67 2.99 2.39 3.06 3.48 2.68
C2RCC_rhow_5x5_depth_in_0_1 1.93 2.57 1.81 1.68 2.09 3.24 1.97 2.14 2.33 1.86
C2X-Complex_rhow_15x15_depth_in_0_1 1.80 2.42 1.71 1.68 1.85 2.43 2.02 1.96 2.28 1.86
C2RCC_rhown_5x5_depth_in_0_1 1.84 2.57 1.73 1.82 1.94 3.05 1.86 2.03 2.31 1.70
C2X-Complex_rhow_5x5_depth_in_0_1 1.73 2.95 2.09 2.05 1.70 4.53 2.20 1.94 2.26 1.72
C2RCC_rhow_15x15_depth_in_0_1 2.05 2.79 1.73 1.73 2.33 2.69 2.07 2.12 2.36 1.99
C2RCC_rhow_9x9_depth_in_0_1 1.88 2.47 1.92 1.98 1.93 2.54 2.19 2.08 2.29 1.79
C2X-Complex_rhown_15x15_depth_in_0_1 2.05 2.47 1.97 1.88 2.23 2.59 2.21 2.22 2.30 2.07
Next in the water column is the depth from 1 to 2 meters, which results are in Tables VI and VII. In this case, datasets processed with C2X-Complex predominated again among the best results. The top performance was achieved with “C2X-Complex_rhow_5x5”, followed by the same configuration with a 9×9 window, which had been the best performer at the 0–1 m depth. Here, the ensemble provided the strongest results, reaching an $R^{2}$ of 0.87 and an RMSE of 1.53 mg/m3, closely followed by CatBoost for both the 5×5 and 9×9 windows.

### v2

@@ 3.3 Buoy data
The data gathered comes from two complementary sources [universidadpolitecnicadecartagenaDatosBoyasUPCT] and [institutomurcianodeinvestigacionydesarrolloagrarioymedioambientalDatosBoyasIMIDA]. The latter is no longer available due to the end of the project related to that website. For simplicity, the initial data source will be referred to as Polytechnic University of Cartagena (UPCT), and the subsequent one as Murcian Institute of Agricultural and Environmental Research and Development (IMIDA). It is hypothesized that both sources are located in the same locations; however, the metadata and information in the data sources do not specify if the buoys are the same. The UPCT dataset encompasses the period from 2017-05-19 to 2024-07-18. Buoys measure at increasing depths, ranging from 0.5 meters to 5.0 meters. In areas where it is feasible to descend to these depths, measurements are typically taken in steps of 0.5 meters. The IMIDA dataset ranges form 2016-06-29 to 2023-08-24 and measurements are taken at surface level (0.0 meters) and then at 1, 2, 3, 4 and 5 meters. In both cases, data is retrieved approximately once a week, despite variations over time and across seasons.
An analysis comparing both datasets at each depth reveals that, in most cases, the differences between the measurements are not significant. Therefore, the merged data goes from 2016-06-29 to 2024-07-18, achieving close to a decade-long dataset.
@@
3.4 Satellite imagery
Sentinel 2 imagery is accessible to the public via the Copernicus Browser [CopernicusBrowser]. Within this interface, users can manually define a bounding box to delineate the area of interest, specify a date range, select the desired level of processing, choose the data source and download SAFE files. These SAFE files correspond to a large area, particularly a tile, and contains an image file for each spectral band, along with metadata. A different approach that is considerably more suitable for the downloading of multiple files is through the Copernicus API, a method that permits the automation of the process. In our case, we use a Python script to achieve this objective.
This script was developed to access the Copernicus Data Space Ecosystem repository through the boto3 [Boto314026Documentation] interface, The dates of interest and the Sentinel 2 tile identifier corresponding to the study area (Tile ID: 30SXG) were specified. The script dynamically constructs the search prefixes based on the date, following the hierarchical structure of the repository. The specific designation for this satellite is "Sentinel 2/MSI/L1C_N0500/<<year>>/<<month>>/<<day>>" or "Sentinel 2/MSI/L1C/<<year>>/<<month>>/<<day>>", depending on the date. For each shot, the product in SAFE format containing the required Tile ID is selected, and all spectral bands are downloaded together with the associated metadata files to a local directory. By downloading L1C products as files in SAFE format, we can process them later on with SNAP.
@@ 4 Methodology
4.1 Buoy Data preprocessing
@@
4.2 Satellite data processing
The SAFE files are processed with SNAP. In order to apply the atmospheric correction algorithms from the C2RCC project the following process is performed:
@@ 4.2 Satellite data processing
    Resample the image to 10 m with the “Nearest” method. The decision of resampling to 10 m rather than downsampling to the native 60 m of certain bands, was driven by the need to maintain high feature granularity. A priori, it is not possible to determine if aggregating spectral information into larger pixels (e.g., 60 m) would improve model performance or, conversely, act as a low-pass filter that obscures spatial variability. By maintaining a 10 m base resolution, the ML models are provided with the highest spatial resolution available, allowing the subsequent window aggregation strategies (from $\times3$ to $\times15$ windows) to systematically explore the optimal scale for noise reduction without sacrificing the underlying spatial structure of the data.
@@
    Crop the subset of the image with the geocoordinates 37.82, -0.867, 37.62, -0.7 for north latitude, west longitude, south latitude, and east latitude bounds, respectively, to delineate the Mar Menor lagoon area for targeted analysis.
@@
    Apply de C2RCC processing for S2-SMI. The outputs selected were TOA reflectantes (uncorrected), normalized water-leaving reflectances (named rhown) and atmospherically corrected angular dependent reflectances (rhow) [SNAPDataProcessors]. While Remote Sensing Reflectance ($R_{rs}$) is also commonly reported in literature, rhow and $R_{rs}$ are linearly related by a factor of $\pi$[petersStepForumC2RCC2020, petersStepForumC2RCC2018], and are more common when dealing with shallow waters [katlaneRemoteSensingTurbidity2024, pereira-sandovalEvaluationAtmosphericCorrection2019]. Therefore, the choice of water-leaving reflectances over $R_{rs}$ does not affect the predictive power or the internal logic of the ML algorithms, as these models are invariant to linear scaling of the input features.
@@
Once the flux is defined, it is possible to automate the process with the SNAP functionality Graph Processing Framework [SNAPGraphProcessing]. This framework is used with XML files that define the processing workflow described above, which is further customized through a properties file to manage the parameters of each step. To fully automate the process, a bash script is then employed. The script iterates through all the dates of interest, applies the workflow specified in the XML file, and saves the resulting outputs as TIFF files. There are three iterations for each date to obtain the image processed with C2RCC, C2X and C2X-Complex.
The TIFF files were processed using Python with the aim of aggregating pixels within windows of different sizes to assess which aggregation scale yields the most reliable results. For each of the C2RCC neural networks, five window sizes were tested: $\times1$, $\times3$, $\times5$, $\times9$, and $\times15$ pixels. In practice, this means that reflectance values at the buoy locations were derived by calculating the mean of the surrounding pixels rather than relying solely on the exact pixel corresponding to the buoy position. This approach was adopted because the processed images often exhibited substantial pixel-to-pixel variation even within very small areas. By averaging across a window, local fluctuations are smoothed, providing a more representative estimate of the reflectance at a given location. The outputs of this procedure were stored as CSV files, with one file generated for each combination of neural network and window size.
4.2.1 Band combinations
Following the ideas from related works previously mentioned, along with more options provided by [delegidoModeloEmpiricoPara2014], many band combinations were explored as inputs for the models. In preliminary tests, formulas from [yeleyMarMenorLagoon2022], [erenaMonitoringCoastalLagoon2019] and [gomez-jakobsenMonitoringChlorophyllConcentration2025] where considered, but were finally discarded due to not improving the models. Therefore, the combinations used are:
@@ 4.2.1 Band combinations
4.3 Datasets
With the buoy and satellite data prepared, the next step is combine them into merged datasets, and obtain a final dataset for each group of depth, processing, and window aggregation. Additionally, C2RCC, C2X and C2X-Complex add another layer of dimensionality because they derived into two versions: atmospherically corrected and normalized reflectances, named rhow y rhown, respectively; whereas TOA reflectances are obtained from one of the previous, and are a single set of bands since they have not undergone any processing. The merging of buoy and satellite data is performed exclusively when both measurement sources are collected on the same date. The proposal in [gomezNewApproachMonitor2021] to include more observations by using satellite images taken one or two days before or after the chlorophyll measurement was considered. However, since chlorophyll is a highly variable parameter, the data matching time window was restricted to a same-day criterion to minimize variance. Sentinel 2 acquisitions over Mar Menor typically occurred between 10:30 and 11:30 [CopernicusBrowser], and buoy measurements are usually taken between 10:00 and 13:00. Consequently, the maximum temporal offset between the two sensing platforms is approximately 2.5 hours. Given the hydrodynamic stability of the lagoon under the selected cloud-free conditions, Chl-a concentrations are not expected to undergo significant biological or spatial fluctuations within this narrow window. This synchronization ensures that the spectral signatures captured by the satellite are representative of the physical samples measured by the in situ sensors, thereby providing a high-fidelity dataset for the training and validation of the ML architectures. At this point, an auxiliary high-chlorophyll class was defined as the 90th percentile of the Chl-a concentration to ensure a balanced representation of scarce high-concentration samples, as well as a categorical variable to include the season of the observation.
@@ 4.3 Datasets
5 Experiments
The experiments conducted are described in this section. The scripts are available in the Github repository https://github.com/Antonio-MI/mar-menor-chl. The experiments were run on four virtual machines with Ubuntu 24.04 LTS, 128 GB of RAM, and an AMD EPYC 9554 64-Core processor each. The code was implemented using Python 3.11.0, and SNAP processing with SNAP 12.0.0.
@@ 5 Experiments
The preliminary training stage began with setting the model parameters to their default values, removing missing entries on the datasets, and training with a 5-fold stratified cross-validation scheme. Using this cross-validation setup (described in more detail in the final training phase), the top ten performing datasets were selected for each processing method. This was based only on the highest R2 and Root Mean Squared Logarithmic Error (RMSLE) values. This means that, across all the algorithms, the best one determines whether a dataset is selected or not.
After preliminary training, hyperparameter optimization was performed using the Optuna framework [akiba2019optuna], which allows to create a “study” instance to optimize an objective function for a particular value specified by the user. Given the high variability of chlorophyll-a concentrations in the Mar Menor, which range from oligotrophic levels to extreme eutrophic peaks, a multi-objective optimization was implemented. The study was configured to simultaneously minimize the Root Mean Squared Logarithmic Error (RMSLE) and maximize the $R^{2}$ score. During this phase, a 4-fold stratified (with the high-chlorophyll class) cross-validation scheme was employed to evaluate those two metrics and tune the models’ hyperparameters. The inclusion of RMSLE as one of the objectives is crucial for this context, as it provides a more balanced penalization of relative errors across different orders of magnitude. On the other hand, the $R^{2}$ accounts for global variance explanation. This process led to the identification of the Pareto front, which consists of the set of optimal solutions where it is impossible to improve one metric without further degrading the other. By exploring this front, Optuna identified hyperparameter configurations that provide a balanced trade-off between overall predictive power and stability across the entire concentration spectrum of the water column.
Given the high dimensionality of the experimental design, which involved 100 independent seeds, multiple water column depths, and various dataset configurations, the total number of required training trials was exceptionally high. To manage this computational load and maximize the system’s efficiency, a parallel execution strategy was implemented. This approach involved parallelizing the Optuna optimization process alongside the execution of individual ML models that support multi-threading.
Once the hyperparameters had been optimized for each dataset–defined by the target Chl-a depth, the processing method, and the reflectance aggregation window–the final training phase was carried out for each depth. Previous iterations of the study revealed that models trained on a single seed could achieve high performance; however, these results exhibited significant sensitivity to the initial data partition, leading to inconsistent estimates of the model’s actual precision. To address this and ensure the statistical robustness of the findings, the entire experimental process was repeated across 100 different random seeds. In the proposed workflow, for each seed, the data were partitioned into a training set (80%) and a test set (20%). This split was stratified using the auxiliary high-chlorophyll class.
Input variables ($X$) were separated from the target ($y$), strictly excluding the high chlorophyll feature from the input to prevent data leakage. For distance-based models (such as KNN or SVR), data were scaled using a RobustScaler before training, and predictions were subsequently back-transformed. In contrast, tree-based models were trained directly on the raw features. The ensemble model was constructed by calculating the simple arithmetic mean of the predictions generated by all individual base models.
Finally, performance metrics ($R^{2}$, RMSLE, and RMSE) were computed for both training and test sets. All results, including best parameters and predictions, were serialized using Pickle to ensure full reproducibility of the experiment. After evaluating all the experiments with different seeds, the results were aggregated to obtain a mean value and its respective standard deviation. Those results are discussed in the following section.
5.1 Results
@@
Model performance was evaluated using three complementary metrics: the root mean squared logarithmic error (RMSLE), the coefficient of determination ($R^{2}$) and the root mean squared error (RMSE). RMSLE was selected because Chl-a data gathered in this study follows a log-normal distribution spanning several orders of magnitude, and therefore the error should be optimized in the logarithmic space. Optimizating this metric would lead to stable predictions if for both low and high Chl-a scenarios. R-Squared, or $R^{2}$, quantifies the proportion of variance in the observed values that is explained by the model, with values closer to 1 indicating a better fit. RMSE measures the average magnitude of the prediction errors, expressed in the same units as the target variable, and lower values denote higher accuracy. Although this metric was not prioritized during model optimization, it is included to maintain comparability with related works.
Formally, given observed values $y_{i}$, predicted values $\hat{y}_{i}$, and their mean $\bar{y}$, the metrics are defined as:
@@
Since models and datasets are grouped for each depth, result are also shown by depth. To ensure conciseness, we only report the test performance metrics of RMSLE and $R^{2}$, but complete test and train metrics, including RMSE as well, are also available in the Supplementary Material.
Metrics shown in the following tables provide the most relevant assessment of the models’ predictive ability. Each table consists of ten rows, one for each of the datasets, determined by type of processing, window aggregation and depth, and a column for each model. Entries on the tables are ordered taking the best value (lower RMSLE or higher R2) of each row, across all the models. The best model for each scenario is highlighted in bold [^3] Linear Regression, ElasticNet and Random Forest were also evaluated, but are omitted here to maintain a legible table while keeping the best-scoring models. Full results are available in the Supplementary Material.. While a robust $R^{2}$ is expected to approach unity, acceptable RMSLE values are inherently context-dependent. In this study, based on the analytical relationship between log-error and variance explanation during training, an RMSLE range between 0.35 and 0.4 is categorized as satisfactory, as it provides a reliable estimation for general monitoring purposes. Consequently, values below 0.35 are considered closer to optimal, whereas values exceeding 0.4 are deemed insufficient for this particular application.
Tables 4 and 5 display the results for the most superficial layer. CatBoost was the best-performing model, obtaining the best metrics for both RMSLE and $R^{2}$, while KNN and SVR also performed consistently. Among datasets, C2RCC and C2X-Complex with and without normalized reflectances, as well as TOA, yielded similar metrics. Likewise, none of the window aggregations performed significantly better than the others, except for the configurations employing $\times1$ or $\times3$ windows, which did not make the top ten results. RMSLE for the best configurations is below 0.34 and $R^{2}$ above 0.75, making predictions for the most superficial depth quite reliable for monitoring.
@@
Table 4: Test RMSLE for depth 0-1 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
TOA_15x15_depth_in_0_1 0.337 ± 0.047 0.362 ± 0.046 0.331 ± 0.049 0.385 ± 0.064 0.409 ± 0.073 0.426 ± 0.038 0.377 ± 0.061
C2X-Complex_rhow_15x15_depth_in_0_1 0.337 ± 0.036 0.356 ± 0.034 0.367 ± 0.042 0.363 ± 0.043 0.392 ± 0.048 0.354 ± 0.033 0.353 ± 0.041
C2RCC_rhow_5x5_depth_in_0_1 0.346 ± 0.035 0.352 ± 0.035 0.356 ± 0.044 0.375 ± 0.039 0.388 ± 0.051 0.368 ± 0.037 0.366 ± 0.039
C2RCC_rhown_5x5_depth_in_0_1 0.347 ± 0.036 0.349 ± 0.034 0.356 ± 0.045 0.363 ± 0.038 0.381 ± 0.047 0.365 ± 0.036 0.355 ± 0.038
C2X-Complex_rhow_9x9_depth_in_0_1 0.349 ± 0.042 0.370 ± 0.042 0.396 ± 0.054 0.378 ± 0.048 0.421 ± 0.053 0.360 ± 0.039 0.363 ± 0.042
C2X-Complex_rhow_5x5_depth_in_0_1 0.364 ± 0.043 0.381 ± 0.048 0.389 ± 0.056 0.409 ± 0.051 0.436 ± 0.050 0.356 ± 0.035 0.396 ± 0.045
C2X-Complex_rhown_15x15_depth_in_0_1 0.359 ± 0.046 0.366 ± 0.041 0.374 ± 0.048 0.383 ± 0.052 0.395 ± 0.043 0.358 ± 0.033 0.377 ± 0.046
C2X-Complex_rhown_9x9_depth_in_0_1 0.362 ± 0.042 0.377 ± 0.046 0.401 ± 0.058 0.386 ± 0.051 0.408 ± 0.058 0.365 ± 0.040 0.382 ± 0.042
C2RCC_rhow_15x15_depth_in_0_1 0.367 ± 0.048 0.375 ± 0.043 0.378 ± 0.062 0.395 ± 0.050 0.416 ± 0.060 0.385 ± 0.045 0.386 ± 0.048
C2RCC_rhow_9x9_depth_in_0_1 0.369 ± 0.041 0.382 ± 0.041 0.384 ± 0.052 0.405 ± 0.045 0.421 ± 0.055 0.382 ± 0.042 0.391 ± 0.043
@@
Table 5: Test $R^{2}$ for depth 0-1 meters. The best model for each scenario is highlighted in bold.
Model CAT ENS KNN LBM MLP SVR XGB
C2RCC_rhow_5x5_depth_in_0_1 0.767 ± 0.082 0.768 ± 0.072 0.756 ± 0.099 0.707 ± 0.118 0.743 ± 0.138 0.706 ± 0.086 0.734 ± 0.100
C2RCC_rhown_5x5_depth_in_0_1 0.763 ± 0.083 0.767 ± 0.070 0.758 ± 0.093 0.719 ± 0.123 0.751 ± 0.113 0.701 ± 0.078 0.742 ± 0.106
TOA_15x15_depth_in_0_1 0.743 ± 0.117 0.725 ± 0.092 0.764 ± 0.131 0.626 ± 0.201 0.659 ± 0.209 0.411 ± 0.091 0.618 ± 0.195
C2X-Complex_rhow_9x9_depth_in_0_1 0.763 ± 0.082 0.736 ± 0.083 0.670 ± 0.122 0.708 ± 0.122 0.656 ± 0.139 0.699 ± 0.098 0.732 ± 0.102
C2X-Complex_rhow_15x15_depth_in_0_1 0.755 ± 0.098 0.729 ± 0.086 0.711 ± 0.110 0.662 ± 0.143 0.704 ± 0.112 0.698 ± 0.111 0.696 ± 0.118
C2X-Complex_rhow_5x5_depth_in_0_1 0.737 ± 0.089 0.709 ± 0.089 0.666 ± 0.126 0.639 ± 0.143 0.647 ± 0.131 0.693 ± 0.107 0.640 ± 0.126
C2RCC_rhow_15x15_depth_in_0_1 0.730 ± 0.098 0.723 ± 0.089 0.733 ± 0.148 0.669 ± 0.132 0.682 ± 0.147 0.677 ± 0.097 0.679 ± 0.131
C2X-Complex_rhown_9x9_depth_in_0_1 0.733 ± 0.092 0.716 ± 0.088 0.671 ± 0.123 0.667 ± 0.128 0.673 ± 0.141 0.689 ± 0.096 0.674 ± 0.109
C2RCC_rhow_9x9_depth_in_0_1 0.726 ± 0.098 0.721 ± 0.086 0.724 ± 0.127 0.658 ± 0.136 0.678 ± 0.158 0.691 ± 0.092 0.691 ± 0.119
C2X-Complex_rhown_15x15_depth_in_0_1 0.709 ± 0.115 0.705 ± 0.100 0.702 ± 0.125 0.639 ± 0.178 0.686 ± 0.113 0.691 ± 0.103 0.640 ± 0.138
The next segment of the water column is between 1 and 2 meters deep. The results for this segment are in Tables 6 and 7. In this case, the datasets that were processed with C2X-Complex predominated among the best results. The top performance was achieved with “C2X-Complex_rhow_15x15”, followed by the same configuration with a $\times9$ window. While SVR also provided the strong results, CatBoost consistently outperformed other models in most scenarios. The metrics for this depth remain robust, with an $R^{2}$ above 0.75 and an RMSLE ranging between 0.35 and 0.4. This performance is interpreted as satisfactory, showing only a marginal increase in error compared to the superficial layer.